# Enhancing Real-World Complex Network Representations with Hyperedge Augmentation

### v1

@@ 4.2.2. Choice of the Feature Extraction Function in HyperAug
GCN+HyperAug(HyperConv) 5.26
GCN+HyperAug(ED-HNN) 3.09
GAT+HyperAug(HyperConv) 5.22
GAT+HyperAug(ED-HNN) 3.04
GraphSAGE+HyperAug(HyperConv) 2.26
GraphSAGE+HyperAug(ED-HNN) 2.74
4.2.3. Identifying the Best Feature Extraction Function for HyperAug
@@
Now that the efficacy of HyperAug has been affirmed through the aforementioned ablation studies, we conduct extensive experiments to identify the optimal feature extraction function for HyperAug. We use GCN, GAT and GraphSAGE as the GNNs for applying HyperAug, and test the performance of HyperAug with either HyperConv or ED-HNN as the feature extraction function, on all 23 node classification datasets we build (described in Section 4.1). The experimental settings for these experiments are set to be the same as the final evaluation of HyperAug, as descrbed in Appendix C.
@@
For clarity, we present in Table 4 the average rankings of those six HyperAug applications, together with the three vanilla GNNs without applying HyperAug. The results clearly show that all HyperAug applications surpass the performance of the vanilla GNNs, which further validate the superior efficacy of HyperAug. Among those HyperAug variants, GraphSAGE with HyperAug(HyperConv) demonstrates to be the most effective combination. In addition, for both GCN and GAT, HyperAug(ED-HNN) outperforms HyperAug(HyperConv), suggesting a positive correlation between the expressiveness of the feature extraction hyper-GNN and the performance gain provided by HyperAug. This is likely due to the enhanced ability of more expressive hyper-GNNs to extract hyperedge information, leading to more informative auxiliary embeddings.
@@
Table 5. Accuracy (%) of HyperAug and corresponding baselines and existing graph augmentation methods on selected datasets. *The results for GCN and GAT on the Cora-CoCitation dataset are directly taken from the GAT paper (Veličković et al., 2018). **The results for HyperConv and ED-HNN on the Cora-CoCitation dataset are directly taken from the ED-HNN paper (Wang et al., 2023). The results clearly show superior performance of HyperAug compared to the baseline GNNs and existing graph augmentation methods.
@@
Method GitHub TwitchDE Cora- CoCitation Brain Lung Cancer Computers
RandomGuess 25.0 50.0 14.3 33.3 33.3 10.0
GCN 87.2 $\pm$ 0.0 65.5 $\pm$ 0.2 81.4 $\pm$ 0.5* 62.5 $\pm$ 0.0 59.6 $\pm$ 0.1 75.6 $\pm$ 4.1
GAT 86.4 $\pm$ 0.1 64.5 $\pm$ 0.4 83.0 $\pm$ 0.7* 62.5 $\pm$ 0.1 59.6 $\pm$ 0.0 74.2 $\pm$ 4.3
GraphSAGE 87.1 $\pm$ 0.2 65.7 $\pm$ 0.1 83.2 $\pm$ 0.1 61.8 $\pm$ 0.2 61.5 $\pm$ 1.5 75.0 $\pm$ 1.9
HyperConv 80.8 $\pm$ 0.1 65.4 $\pm$ 0.2 79.1 $\pm$ 1.0** 62.5 $\pm$ 0.0 59.3 $\pm$ 0.3 84.2 $\pm$ 2.0
ED-HNN 86.2 $\pm$ 0.1 68.1 $\pm$ 0.6 80.3 $\pm$ 1.4** 66.3 $\pm$ 1.3 60.2 $\pm$ 1.4 97.3 $\pm$ 0.2
GCN+DropNode 86.2 $\pm$ 0.1 67.9 $\pm$ 0.7 85.5 $\pm$ 1.0 63.3 $\pm$ 0.2 58.2 $\pm$ 1.3 91.8 $\pm$ 2.1
GCN+DropEdge 86.6 $\pm$ 0.2 67.7 $\pm$ 1.0 86.3 $\pm$ 0.4 63.2 $\pm$ 0.4 60.7 $\pm$ 0.3 92.2 $\pm$ 0.7
GCN+Mixup 85.8 $\pm$ 0.3 67.6 $\pm$ 0.8 85.6 $\pm$ 1.4 65.0 $\pm$ 1.0 59.0 $\pm$ 0.6 87.7 $\pm$ 2.9
GCN+NodeFeatureMasking 85.9 $\pm$ 0.0 67.8 $\pm$ 0.2 85.2 $\pm$ 0.9 64.0 $\pm$ 1.2 59.2 $\pm$ 0.8 91.7 $\pm$ 0.3
GraphSAGE+DropNode 86.2 $\pm$ 0.1 68.3 $\pm$ 0.6 86.4 $\pm$ 0.0 63.3 $\pm$ 0.1 64.5 $\pm$ 0.8 95.9 $\pm$ 0.2
GraphSAGE+DropEdge 86.8 $\pm$ 0.1 67.8 $\pm$ 1.3 87.1 $\pm$ 0.1 64.2 $\pm$ 0.2 64.7 $\pm$ 0.7 95.9 $\pm$ 0.3
GraphSAGE+Mixup 85.9 $\pm$ 0.3 67.1 $\pm$ 0.2 87.2 $\pm$ 1.3 64.2 $\pm$ 1.5 63.4 $\pm$ 0.1 92.2 $\pm$ 0.0
GraphSAGE+NodeFeatureMasking 86.1 $\pm$ 0.0 68.1 $\pm$ 0.7 87.1 $\pm$ 0.5 64.4 $\pm$ 0.4 64.9 $\pm$ 2.3 95.7 $\pm$ 0.2
GCN+HyperAug 87.4 $\pm$ 0.3 67.9 $\pm$ 0.9 86.6 $\pm$ 1.4 66.7 $\pm$ 2.6 63.7 $\pm$ 0.9 98.1 $\pm$ 0.7
GraphSAGE+HyperAug 87.3 $\pm$ 0.2 68.3 $\pm$ 0.3 87.2 $\pm$ 1.8 66.6 $\pm$ 1.6 66.4 $\pm$ 0.2 98.2 $\pm$ 0.7
4.3. Main Results
@@
We thoroughly evaluate the performance of HyperAug against three simple-GNN baselines: GCN, GAT and GraphSAGE, and two hyper-GNN baselines: HyperConv and ED-HNN, on all 23 node classification datasets we build. We include two hyper-GNNs as baselines since these augmented graphs also make use of the hyperedge information. Moreover, we also compare the performance of HyperAug with four different existing graph augmentation methods: DropNode (You et al., 2020), DropEdge (Rong et al., 2020), Mixup (Zhang et al., 2018) and NodeFeatureMasking (Feng et al., 2019a). We evaluate those graph augmentation methods on six selected datasets: MUSAE-GitHub and MUSAE-TwitchDE (featuring virtual hyperedge construction from graph statistics); Cora-CoCitation, GRAND-Brain, and GRAND-LungCancer (featuring virtual hyperedge construction from different data perspectives); and Amazon-Computers (featuring virtual hyperedge construction from different data modalities). Each of dataset chosen in this comparison is the largest graph in their corresponding dataset subgroups. GCN and GraphSAGE are used as the GNNs for the evaluation with the graph augmentation methods. Details about the experimental settings can be found in our Appendix C.
@@
Table 5 summarises the results of the experiments. The results unequivocally demonstrate that HyperAug consistently outperform the vanilla GNN baselines and other graph augmentation methods, across all three types of virtual hyperedge construction strategies. This again highlights the superiority of HyperAug, as well as the idea of incorporating higher-order node relations into graph augmentation methods for real-world complex networks. Besides, HyperAug also outperforms both hyper-GNN baselines, thereby indicating that integrating the original graph with the auxiliary features as augmentations is preferable over only learning from the virtual hyperedges and auxiliary features alone. To fully evaluate the performance of HyperAug, we present in Tables 8, 9, 11, 12, 13, 14 and 10 in Appendix D how it performs with different baselines and different setups.
@@
It can be noted that on the MUSAE-TwitchDE and Cora-Co Citation datasets, although HyperAug is still the best performing graph augmentation method, its performance improvements compared to other graph augmentation methods are not as significant as on other datasets. This can be attributed to the following specific characteristics of the two datasets:
@@
  • The virtual hyperedges for the MUSAE-TwitchDE dataset is constructed by computing from the original graph’s statistics, which does not introduce new information to the graph, thus also limiting HyperAug’s effectiveness in leveraging GNNs’ performance.
@@
  • The virtual hyperedges for the Cora-CoCitation dataset is construct by grouping papers with co-authors, which are not diverse to the co-citation links and may contiain redundant information, compared with the GRAND (gene regulations vs. geometrical information) and Amazon (text vs. image modality) datasets.
@@
These findings also serve as valuable guidance for employing HyperAug in real-world applications, in order to fully release its potential.
4.4. HyperAug with Graph Sampling
Table 6. Statistics of the subgraphs obtained by various samplers, compared with the virtual hypergraph $\mathcal{G}^{\prime}$ produced by HyperAug. GraphSAINT-Edge and GraphSAINT-RW perform the best in preserving HyperAug’s virtual hypergraph characteristics.
MUSAE-GitHub MUSAE-Facebook
Samplers Avg. Node Degree Avg. Hyperedge Degree Clustering Coef. Avg. Node Degree Avg. Hyperedge Degree Clustering Coef.
$\mathcal{G}^{\prime}$ 30.66 0.168 4.590 30.44 0.360 9.905
$\mathcal{G}^{\prime}$ + GraphSAINT-Node 42.59 0.268 13.20 42.72 0.268 13.11
$\mathcal{G}^{\prime}$ + GraphSAINT-Edge 38.63 0.366 11.42 38.66 0.366 11.50
$\mathcal{G}^{\prime}$ + GraphSAINT-RW 35.35 0.203 6.48 35.03 0.206 6.54
$\mathcal{G}^{\prime}$ + RandomNode 2.25 0.020 19.16 2.58 0.037 21.73
$\mathcal{G}^{\prime}$ + RandomHyperedge 72.38 0.282 7.48 72.74 0.279 7.44
Figure 3. Auxiliary feature quality as measured by node classification accuracy of different sampling techniques on the MUSAE Facebook and GitHub datasets. The plots compare the accuracy of the HyperConv model with three variants of GraphSAINT samplers at different subgraph sizes. The size of the subgraph is measured by the number of nodes. The black dashed line indicates the performance of HyperConv trained on the whole graph without any sampling. The accuracy increases as subgraph sizes increased, with both GraphSAINT-RW and GraphSAINT-Node outperforming vanilla HyperConv. GraphSAINT-RW consistently shows superior performance across different sampling sizes.
HyperAug effectively improves the performance of GNNs. However, the creation of auxiliary features relies on the learning process in the virtual hypergraph space, and this could be inefficient with a large number of virtual hyperedges. In this section, we demonstrate how HyperAug augmentation can be integrated with graph sampling to enable scalability on very large graphs. Various hypergraph sampling strategies have been proposed for sampling subgraphs with the purpose of preserving the graph statistics (Choe et al., 2022; Dyer et al., 2021). To perform graph sampling on the virtual hypergraphs, we adopt the same sampling strategies in GraphSAINT (Zeng et al., 2020), but extend them to hyperedges. We propose three graph sampler variants: GraphSAINT-Node (node sampler), GraphSAINT-Edge (edge sampler), and GraphSAINT-RW (random walk sampler).
An intuitive procedure is performed to adopt GraphSAINT for virtual hyperedge sampling: any hyperedges containing at least one node in the sampled subgraph are retained, but all nodes not in the subgraph are masked out from those hyperedges. With this method, we preserve the original virtual hypergraph characteristics in the sampled subgraphs to the maximum extent. We also construct two naïve random samplers as baselines for evaluation: random node sampler and random hyperedge sampler, which randomly sample a subset of nodes or hyperedges from the virtual hypergraph according to a uniform sampling distribution.
We evaluate these samplers on the MUSAE Facebook and GitHub datasets, the two largest graphs among the evaluation datasets. Firstly, we sample subgraphs with various samplers to examine how they preserve the structure of the original graphs. This is measured by three graph statistics: average node and virtual hyperedge degrees, and clustering coefficient. We sample subgraphs multiple times and report the average graph statistics in Table 6. Among all constructed samplers, GraphSAINT-Edge and GraphSAINT-RW perform the best in preserving the graph-level statistics. The two random samplers tend to sample subgraphs with distinct structures from the virtual hypergraphs constructed by HyperAug.
To assess the quality of the auxiliary features generated by HyperAug with various graph samplers, the feature extraction function is fixed to HyperConv. We then directly use the produced auxiliary features to perform node prediction tasks. Figure 3 shows distinct patterns regarding how the accuracy varied with different subgraph sizes. Evaluations were carried out on two distinct datasets, yielding the following key observations:
  • There exists a positive correlation between the size of sampled subgraphs and model accuracy across both datasets.
  • At subgraph sizes of 5000 or larger, both GraphSAINT-RW and GraphSAINT-Node outperform vanilla HyperConv (trained on the whole graph without sampling).
This highlights the potential for improved performance of HyperAug integrated with GraphSAINT samplers. This performance gain can be explained by the reduction of over-smoothing brought by sampling subgraphs. We then apply these three GraphSAINT samplers to all 21 subsequent graphs in the evaluation datasets excluding the citation datasets, extracting subgraphs of 1000, 2000, 3000, and 4000 nodes from each. The results of HyperConv with these three samplers on those graphs are summarised in Figure 4.
Generally, HyperAug with graph sampling techniques enhance HyperConv’s accuracy across most graphs. It is especially significant for node regression datasets like MUSAE-Chameleon, MUSAE-Crocodile, and MUSAE-Squirrel, where the MSE improvement is considerable. Overall, the GraphSAINT samplers, which mostly use simple edge information, can both effectively preserve the virtual hypergraph characteristics, and improve the quality of auxiliary features produced by HyperAug.
5. Conclusion
@@
This work introduces Hyperedge Augmentation (HyperAug), a novel graph augmentation method designed to improve the performance of GNNs in node prediction tasks across diverse real-world large-scale graph datasets. HyperAug enhances the capability of GNNs by generating auxiliary features for each node using virtual hypergraphs. This improvement allows even basic GNN models to surpass the limitations of the 1-WL graph isomorphism test. Empirical results confirm that HyperAug outperforms traditional data augmentation methods in terms of accuracy across various graph types. Additionally, the study showcases HyperAug’s adaptability by integrating it with graph sampling methods, making it applicable to real-world large-scale graphs.
@@
In this study, we demonstrate that augmenting GNNs with a straightforward strategy—concatenating the auxiliary feature derived from hypergraph space with the original feature—significantly boosts their performance. However, several potential methods to leverage the auxiliary feature remain unexplored. These include applying a cross-attention mechanism, as detailed in works (Wei et al., 2020; Vaswani et al., 2017), to effectively integrate the auxiliary and original features. Another unexplored avenue is the concurrent optimisation of the auxiliary and original features using contrastive loss, a technique discussed in ViLT (Kim et al., 2021). Employing these methods has the potential to optimise the integration and utility of the auxiliary features within GNNs, thereby further increasing their overall performance.
Acknowledgements.
@@
This work was performed using the Sulis Tier 2 HPC platform hosted by the Scientific Computing Research Technology Platform at the University of Warwick, the JADE Tier 2 HPC facility, and the Cirrus UK National Tier-2 HPC Service at EPCC. Sulis is funded by EPSRC Grant EP/T022108/1 and the HPC Midlands+ consortium. JADE is funded by EPSRC Grant EP/T022205/1. Cirrus is funded by the University of Edinburgh and EPSRC Grant EP/P020267/1. Zehui Li acknowledges the funding from the UKRI 21EBTA: EB-AI Consortium for Bioengineered Cells & Systems (AI-4-EB) award, Grant BB/W013770/1. Xiangyu Zhao acknowledges the funding from the Imperial College London Electrical and Electronic Engineering PhD Scholarship. Mingzhu Shen acknowledges the funding from the Imperial College London President’s PhD Scholarship. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.

### v2

@@ 4.3.3 Identifying the Best Auxiliary Function
GCN+TopoAug(HyperConv) 5.26
GCN+TopoAug(ED-HNN) 3.09
GAT+TopoAug(HyperConv) 5.22
GAT+TopoAug(ED-HNN) 3.04
GraphSAGE+TopoAug(HyperConv) 2.26
GraphSAGE+TopoAug(ED-HNN) 2.74
@@
Now that the efficacy of TopoAug has been affirmed through the aforementioned ablation studies, we conduct extensive experiments to identify the optimal auxiliary function for TopoAug, denoted as $f_{\text{emb}}$+TopoAug($f_{\text{aux}}$). We use GCN, GAT and GraphSAGE as the embedding GNNs $f_{\text{emb}}$ for applying TopoAug, and test the performance of TopoAug with either HyperConv or ED-HNN as the auxiliary function $f_{\text{aux}}$, on all 23 node classification datasets we build (details of the datasets are described in Section 4.1).
@@
For clarity, we present in Table 4 the average rankings of those six TopoAug applications, together with the three vanilla GNNs without TopoAug. The results clearly show that all TopoAug applications surpass the performance of the vanilla GNNs, which further validate the superior efficacy of TopoAug. Among those TopoAug variants, GraphSAGE with TopoAug(HyperConv) demonstrates to be the most effective combination. In addition, for both GCN and GAT, TopoAug(ED-HNN) outperforms TopoAug(HyperConv), indicating a positive correlation between the expressiveness of the auxiliary hyper-GNN and the performance gain provided by TopoAug. This is likely due to the enhanced ability of more expressive hyper-GNNs to extract hyperedge information, leading to more informative auxiliary features.
@@
Table 5: Accuracy (%) of TopoAug and corresponding baselines and existing graph augmentation methods on selected node classification datasets. *The results for GCN and GAT on the Cora-CoCitation dataset are directly taken from the GAT paper [33]. **The results for HyperConv and ED-HNN on the Cora-CoCitation dataset are directly taken from the ED-HNN paper [34]. The results clearly show superior performance of TopoAug compared to the baseline GNNs and existing graph augmentation methods.
Method Graph Statistics Multi-Perspective Multi-Modality
@@
GitHub TwitchDE Cora- CoCitation Brain Lung Cancer Computers Photos
RandomGuess 25.0 50.0 14.3 33.3 33.3 10.0 10.0
GCN 87.2 $\pm$ 0.0 65.5 $\pm$ 0.2 81.4 $\pm$ 0.5* 62.5 $\pm$ 0.0 59.6 $\pm$ 0.1 75.6 $\pm$ 4.1 29.5 $\pm$ 1.7
GAT 86.4 $\pm$ 0.1 64.5 $\pm$ 0.4 83.0 $\pm$ 0.7* 62.5 $\pm$ 0.1 59.6 $\pm$ 0.0 74.2 $\pm$ 4.3 43.4 $\pm$ 7.4
GraphSAGE 87.1 $\pm$ 0.2 65.7 $\pm$ 0.1 83.2 $\pm$ 0.1 61.8 $\pm$ 0.2 61.5 $\pm$ 1.5 75.0 $\pm$ 1.9 36.6 $\pm$ 6.1
HyperConv 80.8 $\pm$ 0.1 65.4 $\pm$ 0.2 79.1 $\pm$ 1.0** 62.5 $\pm$ 0.0 59.3 $\pm$ 0.3 84.2 $\pm$ 2.0 33.7 $\pm$ 5.9
ED-HNN 86.2 $\pm$ 0.1 68.1 $\pm$ 0.6 80.3 $\pm$ 1.4** 66.3 $\pm$ 1.3 60.2 $\pm$ 1.4 97.3 $\pm$ 0.2 78.6 $\pm$ 1.2
GCN+DropNode 86.2 $\pm$ 0.1 67.9 $\pm$ 0.7 85.5 $\pm$ 1.0 63.3 $\pm$ 0.2 58.2 $\pm$ 1.3 91.8 $\pm$ 2.1 71.1 $\pm$ 0.5
GCN+DropEdge 86.6 $\pm$ 0.2 67.7 $\pm$ 1.0 86.3 $\pm$ 0.4 63.2 $\pm$ 0.4 60.7 $\pm$ 0.3 92.2 $\pm$ 0.7 78.6 $\pm$ 0.6
GCN+Mixup 85.8 $\pm$ 0.3 67.6 $\pm$ 0.8 85.6 $\pm$ 1.4 65.0 $\pm$ 1.0 59.0 $\pm$ 0.6 87.7 $\pm$ 2.9 78.0 $\pm$ 0.1
GCN+NodeFeatureMasking 85.9 $\pm$ 0.0 67.8 $\pm$ 0.2 85.2 $\pm$ 0.9 64.0 $\pm$ 1.2 59.2 $\pm$ 0.8 91.7 $\pm$ 0.3 80.7 $\pm$ 0.6
GraphSAGE+DropNode 86.2 $\pm$ 0.1 68.3 $\pm$ 0.6 86.4 $\pm$ 0.0 63.3 $\pm$ 0.1 64.5 $\pm$ 0.8 95.9 $\pm$ 0.2 69.8 $\pm$ 0.5
GraphSAGE+DropEdge 86.8 $\pm$ 0.1 67.8 $\pm$ 1.3 87.1 $\pm$ 0.1 64.2 $\pm$ 0.2 64.7 $\pm$ 0.7 95.9 $\pm$ 0.3 80.5 $\pm$ 0.6
GraphSAGE+Mixup 85.9 $\pm$ 0.3 67.1 $\pm$ 0.2 87.2 $\pm$ 1.3 64.2 $\pm$ 1.5 63.4 $\pm$ 0.1 92.2 $\pm$ 0.0 71.5 $\pm$ 0.3
GraphSAGE+NodeFeatureMasking 86.1 $\pm$ 0.0 68.1 $\pm$ 0.7 87.1 $\pm$ 0.5 64.4 $\pm$ 0.4 64.9 $\pm$ 2.3 95.7 $\pm$ 0.2 79.3 $\pm$ 0.8
GCN+TopoAug 87.4 $\pm$ 0.3 67.9 $\pm$ 0.9 86.6 $\pm$ 1.4 66.7 $\pm$ 2.6 63.7 $\pm$ 0.9 98.1 $\pm$ 0.7 80.9 $\pm$ 0.3
GraphSAGE+TopoAug 87.3 $\pm$ 0.2 68.3 $\pm$ 0.3 87.2 $\pm$ 1.8 66.6 $\pm$ 1.6 66.4 $\pm$ 0.2 98.2 $\pm$ 0.7 80.9 $\pm$ 0.7
Table 6: MSE ($\downarrow$) of TopoAug and three vanilla GNNs on the node regression datasets (MUSAE-Wiki). The results clearly show superior performacne of TopoAug compared to the vanilla GNN baselines.
Method Squirrel Crocodile Chameleon
GCN 7.319 $\pm$ 0.000 8.761 $\pm$ 0.001 6.779 $\pm$ 0.005
GAT 7.313 $\pm$ 0.007 8.093 $\pm$ 0.054 6.249 $\pm$ 0.261
HyperConv 7.230 $\pm$ 0.002 8.706 $\pm$ 0.000 6.712 $\pm$ 0.001
GCN+TopoAug 6.557 $\pm$ 0.154 4.851 $\pm$ 0.014 5.515 $\pm$ 0.008
GAT+TopoAug 6.049 $\pm$ 0.204 4.875 $\pm$ 0.031 4.665 $\pm$ 0.050
4.4 Main Results
We thoroughly evaluate the performance of TopoAug against three simple-GNN baselines: GCN, GAT and GraphSAGE, and two hyper-GNN baselines: HyperConv and ED-HNN, on all 23 node classification datasets we build. We include the two hyper-GNNs as baselines since the augmented combinatorial complexes constructed by TopoAug also make use of the hyperedge information.
@@ 4.4 Main Results
Moreover, we also compare the performance of TopoAug with four different existing graph augmentation methods: DropNode [40], DropEdge [28], Mixup [42] and NodeFeatureMasking [9]. We evaluate those graph augmentation methods on seven selected datasets: MUSAE-GitHub and MUSAE-TwitchDE (featuring virtual hyperedge construction from graph statistics); Cora-CoCitation, GRAND-Brain, and GRAND-LungCancer (featuring virtual hyperedge construction from different data perspectives); as well as Amazon-Computers and Amazon-Photos (featuring virtual hyperedge construction from different data modalities). Each of dataset chosen in this comparison is among the largest graphs in their corresponding dataset subgroups. GCN and GraphSAGE are used as the GNNs for the evaluation with the graph augmentation methods. Details about the experimental settings and the full results are provided in Appendix C.
@@
Table 5 summarises the results of the experiments. The results unequivocally demonstrate that TopoAug consistently outperforms the vanilla GNN baselines and other graph augmentation methods, across all three types of virtual hyperedge construction strategies. This again highlights the superiority of TopoAug, as well as the idea of incorporating higher-order node relations into graph augmentation methods for real-world complex networks. Besides, TopoAug also outperforms both hyper-GNN baselines, thereby indicating that integrating the original graph with the auxiliary features as augmentations is preferable over only learning from the virtual hyperedges and auxiliary features alone. Among the two backbone embedding GNNs on which TopoAug is applied, GraphSAGE with TopoAug performs better than GCN with TopoAug in general, which is likely due to the more expressive nature of GraphSAGE on those tasks, compared with GCN.
@@
It is noted that on the MUSAE-TwitchDE and Cora-CoCitation datasets, although TopoAug is still the best performing method, its performance improvements compared to other graph augmentation methods are not as significant as on other datasets. This can be attributed to the following specific characteristics of the two datasets:
@@
  • The virtual hyperedges for the MUSAE-TwitchDE dataset are constructed by computing from the original graph’s statistics, which does not introduce new information to the graph, thus also limiting TopoAug’s effectiveness in leveraging GNNs’ performance.
@@
  • The virtual hyperedges for the Cora-CoCitation dataset are constructed by grouping papers with co-authors, which are not diverse to the co-citation links and may contain redundant information, compared with the GRAND (gene regulations vs. geometrical information) and Amazon (text vs. image modality) datasets.
@@
These findings provide valuable guidance for employing TopoAug in real-world applications, in order to fully unleash its potential.
In addition, we also perform supplementary evaluation of TopoAug on the three node regression datasets: MUSAE-Wiki-Chameleon, MUSAE-Wiki-Crocodile and MUSAE-Wiki-Squirrel, against the GCN, GAT and HyperConv baselines, and report their MSEs in Table 6. The results also show consistently and significantly lower MSEs when TopoAug is applied, which further validate the superiority of TopoAug. On those node regression datasets, GAT with TopoAug generally outperforms GCN with TopoAug, since GAT is more expressive than GCN on those tasks. This observation also aligns well with the results on the node classification datasets, which suggests that TopoAug is stable across various types of tasks.
5 Conclusion
@@
This work introduces Topological Augmentation (TopoAug), a novel graph augmentation method designed to improve the performance of GNNs in node prediction tasks across diverse real-world large-scale graph datasets. TopoAug enhances the capability of GNNs by constructing combinatorial complexes from the original graphs using virtual hyperedges, and then generating auxiliary features for each node. This improvement allows even basic GNN models to surpass the limitations of the 1-WL test. Empirical results confirm that TopoAug outperforms traditional graph augmentation methods in terms of node prediction across various data domains.
@@
For future work, there are several potential methods that could be explored to further utilise the auxiliary information. These include applying a cross-attention mechanism, as detailed in works [32, 36], to effectively integrate the auxiliary and original features. Another unexplored avenue is the concurrent optimisation of the auxiliary and original features using contrastive loss, a technique discussed in ViLT [18]. Employing these methods has the potential to optimise the integration and utility of the auxiliary features within GNNs, thereby further increasing their overall performance.
{ack}
@@
This work was performed using the Sulis Tier 2 HPC platform hosted by the Scientific Computing Research Technology Platform at the University of Warwick, the JADE Tier 2 HPC facility, and the Cirrus UK National Tier-2 HPC Service at EPCC. Sulis is funded by EPSRC Grant EP/T022108/1 and the HPC Midlands+ consortium. JADE is funded by EPSRC Grant EP/T022205/1. Cirrus is funded by the University of Edinburgh and EPSRC Grant EP/P020267/1. Xiangyu Zhao acknowledges the funding from the Imperial College London Electrical and Electronic Engineering PhD Scholarship. Zehui Li acknowledges the funding from the UKRI 21EBTA: EB-AI Consortium for Bioengineered Cells & Systems (AI-4-EB) award, Grant BB/W013770/1. Mingzhu Shen acknowledges the funding from the Imperial College London President’s PhD Scholarship. Furthermore, the authors would also like to thank Jing Zeng for her careful proofreading of this manuscript.